Order of stability and bond angles

3 isomers:
(E)-1,2-diisopropylethene
(Z)-1,2-diisopropylethene
1,1-diisopropylethene

What is the order of stability of these three isomers based on their steric energies? and Why?

Also, the order of increasing C=C-C bond angle of these three isomers seem to be: 1,1 isomer < E isomer < Z isomer. Why is that?
Answer #1

also in 1,1 diisopropylethene both alkyl group present in same carbon which experience more force by alkene electron ,cause decrease in angle.
Answer #2

The order of stability based on their steric energies is given below: (E)1,2diisopropylethene<(Z)1,2diisopropylethene<1,1diisopropylethene\left( {\bf{E}} \right){\bf{ - 1,2 - diisopropylethene < }}\left( {\bf{Z}} \right){\bf{ - 1,2 - diisopropylethene < 1,1 - diisopropylethene}} .

The order of increasing C=CC{\bf{C = C - C}} bond angle of three isomers is given below:

1,1diisopropylethene<(E)1,2diisopropylethene<(Z)1,2diisopropylethene{\bf{1,1 - diisopropylethene < }}\left( {\bf{E}} \right){\bf{ - 1,2 - diisopropylethene < }}\left( {\bf{Z}} \right){\bf{ - 1,2 - diisopropylethene}}